# From Rotations to Unitaries: Reversible Quantum Processes and the Emergence of the SU(2)SO(3)SU(2)-SO(3) Isomorphism

### v1

@@ IV From state spaces to physical spaces: $SU(2)\rightarrow SO(3)$
with $\mathcal{U}^{*}\mathcal{U}=\mathds{I}_{\mathds{C}^{2}}$. This restriction of unitarity implies $\det\mathcal{U}=e^{i\lambda}$, $\lambda\in\mathds{R}$. Imposing a group structure to (15), we demand $\det\mathcal{U}=1$. This way, two (or more) successive transformations is still unitary and with unit determinant. Hence, $\mathcal{U}\in SU(2)$ and we parametrize the group elements according to the section II: $\mathcal{U}(\hat{n},\alpha)=e^{-i\frac{\alpha}{2}\hat{n}\cdot\vec{\sigma}}$. The transition from (8) to (10) can be reverted. Thus, $\mathcal{U}(\hat{n},\alpha)$ induces the rotation $\mathcal{R}(\hat{n},\alpha)$ given by (5). We are finally led to $\phi^{-1}(\mathcal{U}(\hat{n},\alpha))=\mathcal{R}(\hat{n},\alpha)$.
@@
We begin by pointing out the isomorphism between the Lie algebra $su(2)$ of the group $SU(2)$ and the physical space $\mathds{D}$. Explicitly,
$su(2)=\{u,\mbox{ 2$\times$2 complex matrices, }u^{*}=-u;\,[\ ,\,]\}.$ (16)
As usual, the commutator reads $u_{1},u_{2}]=u_{1}u_{2}-u_{2}u_{1}$, for arbitrary elements $u_{1}$ and $u_{2}$ in the algebra.
@@
which makes explicit the isomorphism $su(2)\cong\mathds{D}$,
@@
[…]
@@
$\mathcal{R}_{kj}=\frac{1}{2}Tr(\mathcal{U}\sigma_{j}\,\mathcal{U}^{*}\sigma_{k})$ (25)
@@
Figure 2: Picture of $SU(2)$ and $SO(3)$ via the isomorphism $su(2)\cong\mathds{D}$.
Eq. (25) shows that both $\mathcal{U}$ and $\mathcal{U}$ are mapped to the same element in $SO(3)$, resulting in the two-to-one (quasi-)isomorphism, as previously stated. Moreover, it is instructive to show that $\mathcal{R}_{kj}$ is indeed an orthogonal matrix. For that, we express (22) in terms of the basis elements $\sigma_{i}$’s,
@@ V Conclusion and Perspectives
In this work, we have explored a concrete physical system to operationally reconstruct the well-known (quasi-) isomorphism between the groups $SU(2)$ and $SO(3)$. Rather than treating this correspondence as a purely mathematical artifact, we built it from first principles using physical procedures such as state preparation, reversible quantum evolutions, and the structure of CPTP maps. This allowed us to realize the isomorphism through a sequence of physically meaningful and experimentally motivated steps, firmly rooted in the geometry of the Bloch sphere and the algebra of quantum operations.
Our construction not only reinforces the deep interplay between group theory and quantum mechanics but also offers a clear operational standpoint through which these connections emerge. By anchoring the abstract isomorphism in transformations that can be implemented in practice—rotations in physical space and unitaries in Hilbert space—we provide a concrete interpretation that enriches the usual formal presentation.
@@
$\rho=\frac{1}{4}\left(S_{00}+\sum_{i}r_{i}S_{i0}+\sum_{j}S_{0j}+\sum_{i,j}t_{ij}S_{ij}\right),$
@@ VII Appendix
Let $\Phi:L(\mathds{C}^{n})\rightarrow L(\mathds{C}^{n})$ be a CPTP map. Suppose that $\Phi$ admits an inverse $\Phi^{-1}$ which is also CPTP. Then there exists a unitary $\mathcal{U}$ on $\mathds{C}^{2}$ such that $\Phi(\cdot)=\mathcal{U}(\cdot)\,\mathcal{U}^{*}$ and $\Phi^{-1}(\cdot)=\mathcal{U}^{*}(\cdot)\,\mathcal{U}$. Conversely, if $\Phi$ is a CPTP map whose Kraus representation has only one operator, say $\Phi(\cdot)=\mathcal{U}(\cdot)\,\mathcal{U}^{*}$, then $\Phi$ is invertible, $\Phi^{-1}(\cdot)=\mathcal{U}^{*}(\cdot)\,\mathcal{U}$ and $\Phi^{-1}$ is also CPTP.

### v2

@@ IV From state spaces to physical spaces: $SU(2)\rightarrow SO(3)$
with $\mathcal{U}^{*}\mathcal{U}=\mathds{I}_{\mathds{C}^{2}}$. This restriction of unitarity implies $\det\mathcal{U}=e^{i\lambda}$, $\lambda\in\mathds{R}$. Because we are aiming the group $SU(2)$, we can consider only unitary matrices with determinant equals one. In fact, if $\mathcal{U}$ is such that $\det\mathcal{U}=e^{i\lambda}$, define
$\mathcal{W}:=\frac{\mathcal{U}}{\sqrt{\det\mathcal{U}}}.$
Therefore, not only $\det\mathcal{W}=1$ but also $\mathcal{W}$ defines the same channel as $\mathcal{U}$,
$\mathcal{W}\rho\mathcal{W^{*}}=\frac{\mathcal{U}}{\sqrt{\det\mathcal{U}}}\rho\frac{\mathcal{U^{*}}}{\sqrt{\overline{\det\mathcal{U}}}}=\frac{1}{|\det\mathcal{U}|^{2}}\mathcal{U}\rho\mathcal{U}^{*}=\mathcal{U}\rho\mathcal{U}^{*}$ (15)
Hence, with no loss of generality, it is enough to consider only $\mathcal{U}\in SU(2)$ and we parametrize the group elements according to the section II: $\mathcal{U}(\hat{n},\alpha)=e^{-i\frac{\alpha}{2}\hat{n}\cdot\vec{\sigma}}$. The transition from (5) to (8) can be reverted. That is, starting from the right hand side of (7), all the trigonometrical identities can be used in the opposite order of the ones taken in the steps from $SO(3)$ to $SU(2)$. Thus, $\mathcal{U}(\hat{n},\alpha)$ induces the rotation $\mathcal{R}(\hat{n},\alpha)$ given by (4). We are finally led to $\phi^{-1}(\mathcal{U}(\hat{n},\alpha))=\mathcal{R}(\hat{n},\alpha)$.
@@
We begin by pointing out the bijection between the Lie algebra $su(2)$ of the group $SU(2)$ and the physical space $\mathds{D}$. Explicitly, $su(2)$ is the set of complex traceless matrices $u$ such that $u^{*}=-u$, equipped with the commutator $u_{1},u_{2}]=u_{1}u_{2}-u_{2}u_{1}$, for arbitrary elements $u_{1}$ and $u_{2}$ in the algebra.
@@
which makes explicit the one-to-one so stated relation $su(2)\cong\mathds{D}$,
@@
where $u\mathbin{{\mathop{\longmapsto}\limits^{T_{\mathcal{U}}}}}\mathcal{U}u\,\mathcal{U}^{*}$. We observe that $T_{\mathcal{U}}$ is linear and for $\mathcal{U},\mathcal{V}\in SU(2)$
@@
$\mathcal{R}_{kj}=\frac{1}{2}Tr(\mathcal{U}\sigma_{j}\,\mathcal{U}^{*}\sigma_{k}).$ (24)
@@
Figure 2: Picture of $SU(2)$ and $SO(3)$ via the $su(2)\cong\mathds{D}$ bijection.
Eq. (24) shows that both $\mathcal{U}$ and $\mathcal{U}$ are mapped to the same element in $SO(3)$, resulting in the two-to-one homomorphism, as previously stated.
The computations above explores the bijection between two (apparently) distinct vector spaces to obtain the advocated homomorphism. One could, however, reach the same result with only a group theoretical standpoint. To see this, firstly we point out that Eq. (9) indicates that
$-\mathcal{U}(\hat{n},\alpha)$ $=-\cos\frac{\alpha}{2}\mathds{I}-i\sin\frac{\alpha}{2}(-\hat{n}\cdot\vec{\sigma})$
$=\cos\left(\frac{2\pi-\alpha}{2}\right)\mathds{I}-i\sin\left(\frac{2\pi-\alpha}{2}\right)(-\hat{n}\cdot\vec{\sigma})$
$=\mathcal{U}(-\hat{n},2\pi-\alpha).$
On the other hand,
$\mathcal{R}(-\hat{n},2\pi-\alpha)=\mathcal{R}(\hat{n},\alpha),$ (25)
which can be concluded by geometric grounds: these formulae shows that rotating around the opposite axis $\hat{n}$ produces the same effect as rotating around the original axis $\hat{n}$, but at the angle $2\pi-\alpha$. This simple observation also makes it clear that two distinct unitaries, namely $\mathcal{U}$ and $\mathcal{U}$ in $SU(2)$, are mapped to the same rotation in $SO(3)$.
It is instructive to show that $\mathcal{R}_{kj}$ is indeed an orthogonal matrix. For that, we express (21) in terms of the basis elements $\sigma_{i}$’s,
@@
Although one can argue that the correspondence between unitaries and rotations is only a two-to-one homomorphism, the situation is indeed a bijection when we restrict our discussion to invertible quantum channels acting on a qubit. Because these maps are of the form
$\Phi(\cdot)=\mathcal{U}\ (\cdot)\,\mathcal{U}^{*},$ (29)
with $\mathcal{U}\in SU(2)$, we note that replacing $\mathcal{U}$ by $\mathcal{U}$ yields the same quantum channel. Therefore, what emerges is in fact an isomorphism between the group of reversible quantum maps and rotations in $\mathds{R}^{3}$.
We conclute this section with an alternative route that could have been taken. Our construction could equivalently begin from the direction $SU(2)\rightarrow SO(3)$. This alternative route is indeed very natural: the role of unitaries is already justified by the discussion of reversible CPTP maps, and such unitaries induce rotations on the Bloch sphere in a straightforward manner. Nevertheless, we adopt the opposite order (starting from transformations of physical space and then considering their induced action on the state space) because our treatment is grounded in an operational perspective. In practice, one first specifies a preparation by selecting a physical direction, and only afterwards considers how transformations act on these preparations. Moreover, both in physical space and in the space of quantum states, the set of mathematically admissible isometries is larger than the set of physically implementable transformations. To establish a meaningful group-theoretic connection between the two domains, it is therefore necessary to restrict attention to the class of transformations that correspond operationally on both sides—namely, proper rotations in $\mathds{R}^{3}$ and their unitary counterparts in $SU(2)$.
@@ V Conclusion and Perspectives
In this work, we have explored a concrete physical system to operationally reconstruct the well-known two-to-one homomorphism between the groups $SU(2)$ and $SO(3)$. Rather than treating this correspondence as a purely mathematical artifact, we built it from first principles using physical procedures such as state preparation, reversible quantum evolutions, and the structure of CPTP maps. This allowed us to realize the homomorphism through a sequence of physically meaningful and experimentally motivated steps, firmly rooted in the geometry of the Bloch sphere and the algebra of quantum operations.
Our construction not only reinforces the deep interplay between group theory and quantum mechanics but also offers a clear operational standpoint through which these connections emerge. By anchoring the abstract homomorphism in transformations that can be implemented in practice—rotations in physical space and unitaries in Hilbert space—we provide a concrete interpretation that enriches the usual formal presentation.
@@
$\rho=\frac{1}{4}\left(S_{00}+\sum_{i}r_{i}S_{i0}+\sum_{j}s_{j}S_{0j}+\sum_{i,j}t_{ij}S_{ij}\right),$
@@ VII Appendix
Let $\Phi:L(\mathds{C}^{n})\rightarrow L(\mathds{C}^{n})$ be a CPTP map. Suppose that $\Phi$ admits an inverse $\Phi^{-1}$ which is also CPTP. Then there exists a unitary $\mathcal{U}$ on $\mathds{C}^{n}$ such that $\Phi(\cdot)=\mathcal{U}(\cdot)\,\mathcal{U}^{*}$ and $\Phi^{-1}(\cdot)=\mathcal{U}^{*}(\cdot)\,\mathcal{U}$. Conversely, if $\Phi$ is a CPTP map whose Kraus representation has only one operator, say $\Phi(\cdot)=\mathcal{U}(\cdot)\,\mathcal{U}^{*}$, then $\Phi$ is invertible, $\Phi^{-1}(\cdot)=\mathcal{U}^{*}(\cdot)\,\mathcal{U}$ and $\Phi^{-1}$ is also CPTP.